# Inferring activity from the flow field around active colloidal particles using deep learning

### v1

@@ Appendix A Solving the Stokes equation to compute flow around active particles
The above integrals can be solved by expanding the boundary fields in TSH (tensorial spherical harmonics) (Hess, 2015). The TSH are defined as $Y_{\alpha_{1}\dots\alpha_{l}}^{(l)}(\hat{\mathbf{b}})=(-1)^{l}\,b^{(l+1)}\,\nabla_{\alpha_{1}}\dots\nabla_{\alpha_{l}}\frac{1}{b}$. The expansion of the single layer density $\mathbf{q}$ in TSH is:
$\mathbf{q}(\mathbf{b})$ $=\sum_{l=1}^{\infty}w_{l}\,\mathbf{Q}^{(l)}\cdot\mathbf{Y}^{(l-1)}(\mathbf{\hat{\mathbf{b}}}),\qquad\mathbf{Q}^{(l)}=\tilde{w}_{l}\int\mathbf{q}\,\mathbf{Y}^{(l-1)}(\mathbf{\hat{\mathbf{b}}}),$ (9a)
Here, $\mathbf{b}$ is the radius vector of active colloidal sphere, while the $l-$dependent constants are: $w_{l}=\frac{1}{(l-1)!(2l-3)!!}$ and ${w}_{l}=\frac{2l-1}{4\pi b^{2}}$. Here, we have defined $M!!=M(M-2)(M-4)\dots$, while $\mathbf{\Delta}^{(l)}$ is a tensor of rank $2l$ projecting the symmetric irreducible part of a tensor of rank $l$ (Hess, 2015). In the above, and throughout the paper, a maximal contraction of two tensors is denoted by a dot product.
The boundary integral of Eq.(8) can be solved exactly as follows: (a) expanding the boundary fields in TSH, (b) performing a Taylor expansion of kernels (the Green’s function) about the center of the particles, and (c) using the orthogonality of TSH (Singh et al., 2015; Singh, 2018). The final result is the following summation:
@@
Thus, the flow due to $l$th mode decays as $r^{-l}$, since $\mathbf{G}(\mathbf{r})\propto r^{-1}$.
The equation (1) in the Fourier space is:
@@
We define Fourier transform of a function $\varphi(\mathbf{r})$ and Dirac delta function $\delta(\mathbf{r})$ as
[…]
@@
The coefficients of single layer density $\mathbf{Q}^{(l)}$ are $l$-th rank tensors, symmetric irreducible in their last $l-1$ indices. The reducible tensors $\mathbf{Q}^{(l)}$ can be written in irreducible forms as: $\mathbf{Q}^{(l\sigma)}=\mathbf{P}^{(l\sigma)}\cdot\mathbf{Q}^{(l)}$ Turk et al. (2024). In the above expression, the operator $\mathbf{P}^{(l\sigma)}$ projects symmetric irreducible ($\mathbf{P}^{(ls)}=\mathbf{\Delta}^{(l)}$), antisymmetric ($\mathbf{P}^{(la)}=\mathbf{\Delta}^{(l-1)}\mathbf{\mathbf{\varepsilon}}$), and trace ($\mathbf{P}^{(lt)}=\mathbf{\delta}$) parts of a reducible tensor of rank-$l$ (Hess, 2015). Here $\mathbf{\Delta}^{(l)}$ is a tensor of rank $2l$ projecting the symmetry irreducible part of a tensor of rank $l$, $\mathbf{\varepsilon}$ is the Levi-Civita tensor and $\mathbf{\delta}$ is the Kronecker delta (Hess, 2015). $\mathbf{Q}^{(l\sigma)}$ are irreducible tensors of rank $l$, $l-1$ and $l-2$ for $\sigma=s,a,$ and $t$ respectively. The final expression of the flow in terms of irreducible components can be written as per Eq.(2) of the main text. For simplicity and less error in computation the symmetric-dipole ($2s$) can also be written as two forces acting along $\mathbf{p}$ and $-\mathbf{p}$ each $\Delta a$ length apart from the center of particle. This formulation of $2s$ doesn’t keep the traceless property, but that is taken care in the flow field by the incompressibility condition.

### v3

@@ Appendix A Solving the Stokes equation to compute flow around active particles
$\mathbf{v}(\mathbf{r})=-\int\mathbf{G}(\mathbf{r}-\mathbf{r}^{\prime})\cdot\mathbf{q}(\mathbf{r}^{\prime})dS$ (8)
@@
[…]
Here, the $l$-dependent constants are: ${w}_{l}=\frac{2l-1}{4\pi b^{2}}$ and $\tilde{w}_{l}=\frac{1}{(l-1)!(2l-3)!!}$ (Hess, 2015). In the above, and throughout the paper, a maximal contraction of two tensors is denoted by a dot product. The boundary integral of Eq.(8) can be solved exactly as follows: (a) expanding the boundary fields in TSH, (b) performing a Taylor expansion of kernels (the Green’s function) about the center of the particles, and (c) using the orthogonality of TSH (Singh et al., 2015; Singh, 2018). The final result is the following summation:
@@
Thus, the flow due to $l$th mode decays as $\nabla^{l-1}\mathbf{G}(\mathbf{r})$. In three-dimensions $\mathbf{G}(\mathbf{r})\propto r^{-1}$, and thus, the flow due to any of the $l\sigma$ mode as $r^{-l}$ (Singh et al., 2015; Deb & Singh, 2025), where $r$ is the distance from the source. The equation (1) in the Fourier space is:
@@
Here, $\mathbf{I}$ is the identity tensor and $\mathbf{k}$ is Fourier wave vector. We define Fourier transform of a function $\varphi(\mathbf{r})$ and Dirac delta function $\delta(\mathbf{r})$ as
${\varphi}(\mathbf{k})=\int\varphi(\mathbf{r})\,e^{-i\mathbf{k}\cdot\mathbf{r}}\,d\mathbf{r},\qquad\delta(\mathbf{r})=\frac{1}{(2\pi)^{3}}\int e^{i\mathbf{k}\cdot\mathbf{r}}\,d\mathbf{k}.$ (12)
@@
The coefficients of single layer density $\mathbf{Q}^{(l)}$ are $l$-th rank tensors, symmetric irreducible in their last l1l-1 indices. The reducible tensors $\mathbf{Q}^{(l)}$ can be written in irreducible forms as: $\mathbf{Q}^{(l\sigma)}$, where $\sigma=s$ represents symmetric-traceless part of the tensor, while $\sigma=a$ represents the antisymmetric part, and $\sigma=t$ denotes the trace. Here, $\mathbf{Q}^{(l\sigma)}$ are irreducible tensors of rank $l$, l1l-1 and l2l-2 for $\sigma=s,a,$ and $t$ respectively. The final expression of the flow - Eq.(2) of the main text - in terms of irreducible components is obtained when Eq.(15) is written as a sum over the irreducible components of $\mathbf{Q}^{(l)}$.
Appendix B Quantitative Analysis
B.1 Root Mean Square Error
The convergence plot (Average of Mean Squared Error over batches of 32 vs Epoch) is given in Fig.5 of the main text. It shows the good fit in neural network. The network was tested over 600 test cases in each of the single particle, multiple particle and variable particle instances. The RMSE(Root Mean Squared Error) of the true force density versus the one predicted by the network over the cases is plotted for each of the instances as histogram Fig.6. The RMSE is calculated over all grid points for a single instance.
B.2 Variable Resolution
In Fig.7, we demonstrate that our network works for variable spatial resolution. In the main text, we have set the spatial resolution to be $128\times 128$ while we vary number of particles. Here, we plot the loss function as a function of epochs for two different resolutions ($100\times 100$ and $200\times 200$) for an inference study of three particles.
Figure 6: Histogram of root mean squared error from the predictions by CNN.
Figure 7: Loss function as a function of epochs for two different resolutions of the data.
B.3 Ablation Study
The advantages of CNN in this particular problem is studied by checking the solutions without the components of CNN. To evaluate it, the dataset was trained with just convolution layers without any activation functions, which essentially makes it a linear map. The comparison of this case against the CNN with activation function is shown by the convergence plot in Fig.8. The former is not converging with similar iterations where as the later converges rapidly. CNNs learn non-linear mappings using activation functions. Thus, we demonstrate the importance of non-linear mappings in deep learning by showing that CNN with activation function is essential for the model presented here to be effectively trained against data.
Figure 8: Ablation study. Measuring loss with epochs using CNN with activation functions (LEFT) and without activation function (RIGHT).